# Sassha: Sharpness-aware Adaptive Second-order Optimization with Stable Hessian Approximation

### v1

@@ 5.2 Language Modeling
Recent studies have shown the potential of second-order methods for pretraining language models. Here, we first evaluate how Sassha performs on this task. Specifically, we train GPT1-mini, a scaled-down variant of GPT1 (Radford et al., 2019), on Wikitext-2 dataset (Merity et al., 2022) using various methods including Sassha and compare their results (see the left of Table 3). Our results show that Sassha achieves the lowest perplexity among all methods including Sophia-H (Liu et al., 2024), a recent method that is designed specifically for language modeling tasks and sets state of the art, which highlights generality in addition to the numerical advantage of Sassha.
@@ 5.3 Comparison to SAM
We posit that this is potentially due to the robustness of Sassha to the block heterogeneity inherent in Transformer architectures, where the Hessian spectrum varies significantly across different blocks. This characteristic is known to make SGD perform worse than adaptive methods like Adam on Transformer-based models (Zhang et al., 2024). Since Sassha leverages second-order information via preconditioning gradients, it has the potential to address the ill-conditioned nature of Transformers more effectively than SAM with first-order methods.
To push further, we conducted additional experiments. First, we allocate more training budgets to SAM to see whether it compares to Sassha. The results are presented in Table 4. We find that SAM still underperforms Sassha, even though it is given more budgets of training iterations over data or wall-clock time. Furthermore, we also compare Sassha to more advanced variants of SAM including ASAM (Kwon et al., 2021) and GSAM (Zhuang et al., 2022), showing that Sassha performs competitively even to these methods (Appendix I). Notably, however, these variants of SAM require a lot more hyperparameter tuning to be compared.
6 Further Analysis
@@ 6.1 Robustness
6.2 Stability
It turns out that this problem stems from the preconditioning matrix $D$ being too small; i.e., the distribution of diagonal entries in the preconditioning matrix gradually shifts toward zero values (Figure 4(c)); as a result, $D^{-1}$ becomes too large, creating large steps. This progressive increase in near-zero diagonal Hessian entries is precisely due to the sharpness minimization scheme that we introduced; it penalizes the Hessian eigenspectrum to yield flat solutions, yet it could also make training unstable if taken naively. By including square-root, the preconditioner are less situated near zero, effectively suppressing the risk of large updates, thereby stabilizing the training process. We validate this further by showing its superiority to other alternatives including damping and clipping in Section E.1.
We also provide an ablation analysis for the absolute-value function in Section E.2, which demonstrates that it increases the stability of Sassha in tandem with square-root.
(a) Lazy Hessian
(b) $\widehat{H}$ change
(c) Local sensitivity
@@ 6.4 Cost
Furthermore, we can leverage a momentum of gradients in the perturbation step to reduce the cost. This variant M-Sassha requires only $1.3$GCs with minimal decrease in performance. Notably, M-Sassha still outperforms standard first-order methods like SGD and AdamW (Appendix K).
@@ 7 Conclusion
In this work, we focus on addressing the issue of poor generalization in approximate second-order methods. To this end, we propose a new method called Sassha that stably minimizes sharpness within the framework of second-order optimization. Sassha converges to flat solutions and achieves state-of-the-art performance within this class. Sassha also performs competitively to widely-used first-order, adaptive, and sharpness-aware methods. Sassha achieves this in robust, stable, and efficient ways without incurring much cost or requiring extra hyperparameter tuning. All of these are rigorously assessed with extensive experiments.
Nonetheless, there are many limitations to be addressed in this work for more improvements which may include, but are not limited to, extending experiments to extreme scales of various models and different data, and consolidating theoretical foundations. Seeing it as an exciting opportunity, we plan to investigate further in future work.
8 Acknowledgement
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (IITP-2019-0-01906, Artificial Intelligence Graduate School Program (POSTECH), RS-2022-II220959, (part2) Few-Shot learning of Causal Inference in Vision and Language for Decision Making, and RS-2024-00338140, Development of learning and utilization technology to reflect sustainability of generative language models and up-to-dateness over time), and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) ( NRF-2022R1F1A1064569, RS-2023-00210466, RS-2023-00265444).
@@ Appendix A Sharpness Measurements for Other Settings
Table 7: Sharpness measurements of the solutions found by seven different optimizers and their generalization on CIFAR-10/100 and Wikitext-2. Approximate second-order methods tend to yield highly sharp solutions and poor generalization compared to SGD; Sassha and M-Sassha effectively recover this. Here, we measure sharpness in terms of maximum Hessian eigenvalue $\lambda_{max}(H)$, trace of Hessian $\operatorname{tr}(H)$, worst-case sharpness $\delta L_{\text{worst}}$, and average sharpness $\delta L_{\text{avg}}$, along with generalization using validation loss $L_{val}$ and accuracy $\text{Acc}_{val}$.
@@
$\lambda_{max}(H)$ $\operatorname{tr}(H)_{\times 10^{3}}$ $\delta L_{\text{worst}}$ $\delta L_{\text{avg}\times 10^{-3}}$ $L_{\text{val}}$ $\text{Acc}_{\text{val}}$
@@
ResNet20 SGD $107_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 4.37}}$ $1.38_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.01}}$ $0.840_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.304}}$ $0.690_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.39}}$ $0.295_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.008}}$ $92.03_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.32}}$
SAM $58_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 2.98}}$ $0.73_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.04}}$ $0.171_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.038}}$ $0.461_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.24}}$ $0.119_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.002}}$ $92.85_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.07}}$
Sophia-H $3606_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 303}}$ $31.24_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 2.628}}$ $6.120_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.634}}$ $18.11_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1}}$ $0.316_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.002}}$ $91.81_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.27}}$
AdaHessian $23048_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 29932}}$ $189.48_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 240.55}}$ $4.538_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.634}}$ $198.66_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 266}}$ $0.260_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.006}}$ $92.00_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.17}}$
Shampoo $647066_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 419964}}$ $3899.5_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1825}}$ $166.3_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 48.00}}$ $2177189_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1628993}}$ $0.381_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.028}}$ $88.55_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.83}}$
M-Sassha $129_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 17}}$ $1.58_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.08}}$ $1.551_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.684}}$ $1.025_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.36}}$ $0.234_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.003}}$ $92.36_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.23}}$
Sassha $78_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 5.09}}$ $0.86_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.03}}$ $0.184_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.053}}$ $0.388_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.704}}$ $0.209_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.001}}$ $92.98_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.05}}$
ResNet32 SGD $56_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 5.10}}$ $0.80_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.04}}$ $0.560_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.219}}$ $0.196_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.146}}$ $0.309_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.002}}$ $92.69_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.06}}$
SAM $45_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 2.67}}$ $0.58_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.02}}$ $0.107_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.005}}$ $0.753_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.351}}$ $0.128_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.001}}$ $93.89_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.13}}$
Sophia-H $7167_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 2755}}$ $18.82_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 5.50}}$ $9.399_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 2.283}}$ $7.915_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 3.397}}$ $0.394_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.010}}$ $91.99_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.08}}$
AdaHessian $1746_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1018}}$ $17.06_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 10.24}}$ $4.599_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.710}}$ $5.518_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 3.623}}$ $0.278_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.006}}$ $92.48_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.15}}$
Shampoo $717553_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 93129}}$ $4523_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 629.7}}$ $162.1_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 123.2}}$ $105322_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 82246}}$ $0.348_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.008}}$ $90.23_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.24}}$
M-Sassha $283_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 10}}$ $3.96_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.10}}$ $2.986_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.133}}$ $1.300_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.969}}$ $0.211_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.010}}$ $93.18_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.30}}$
Sassha $47_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 1.88}}$ $0.59_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.02}}$ $0.136_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.019}}$ $0.714_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.090}}$ $0.177_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.002}}$ $94.09_{{\color[rgb]{0.4,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.4,0.4,0.4}\pgfsys@color@gray@stroke{0.4}\pgfsys@color@gray@fill{0.4}\pm 0.24}}$

### v2

@@ 5.2 Language Modeling
Recent studies have shown the potential of second-order methods for pretraining language models. Here, we first evaluate how Sassha performs on this task. Specifically, we train GPT1-mini, a scaled-down variant of GPT1 (Radford et al., 2019), on Wikitext-2 dataset (Merity et al., 2022) using various methods including Sassha and compare their results (see the left of Table 3). Our results show that Sassha achieves the lowest perplexity among all methods including Sophia-H (Liu et al., 2024), a recent method that is designed specifically for language modeling tasks and sets state of the art, which highlights generality in addition to the numerical advantage of Sassha. We further evaluate Sassha for GPT2 and provide results in Appendix J.
@@ 5.3 Comparison to SAM
To further investigate these findings, we conducted additional experiments. First, we allocate more training budgets to SAM to see whether it compares to Sassha. The results are presented in Table 4. We find that SAM still underperforms Sassha, even though it is given more budgets of training iterations over data or wall-clock time. Furthermore, we also compare Sassha to more advanced variants of SAM including ASAM (Kwon et al., 2021) and GSAM (Zhuang et al., 2022), showing that Sassha performs competitively even to these methods (Appendix K). Notably, however, these variants of SAM require a lot more hyperparameter tuning to be compared.
We suspect that this may be due to the robustness of Sassha to the block heterogeneity inherent in Transformer architectures, where the Hessian spectrum varies significantly across different blocks. This characteristic is known to make SGD perform worse than adaptive methods like Adam on Transformer-based models (Zhang et al., 2024). Since Sassha leverages second-order information via preconditioning gradients, it has the potential to address the ill-conditioned nature of Transformers more effectively than SAM with first-order methods.
@@ 6.1 Robustness
6.2 Stability
It turns out that this problem stems from the preconditioning matrix $D$ being too small; i.e., the distribution of diagonal entries in the preconditioning matrix gradually shifts toward zero values (Figure 4(c)); as a result, $D^{-1}$ becomes too large, creating large steps. This progressive increase in near-zero diagonal Hessian entries is precisely due to the sharpness minimization scheme that we introduced; it penalizes the Hessian eigenspectrum to yield flat solutions, yet it could also make training unstable if taken naively. By including square-root, the preconditioner are less situated near zero, effectively suppressing the risk of large updates, thereby stabilizing the training process. We validate this further by showing its superiority to other alternatives including damping and clipping in Section F.1.
We also provide an ablation analysis for the absolute-value function in Section F.2, which demonstrates that it increases the stability of Sassha in tandem with square-root.
@@ 6.3 Efficiency
(a) Lazy Hessian
(b) $\widehat{H}$ change
(c) Local sensitivity
@@ 6.4 Cost
Furthermore, we can leverage a momentum of gradients in the perturbation step to reduce the cost. This variant M-Sassha requires only $1.3$GCs with minimal decrease in performance. Notably, M-Sassha still outperforms standard first-order methods like SGD and AdamW (Appendix M).
@@ 7 Conclusion
In this work, we focus on addressing the issue of poor generalization in approximate second-order methods. Our empirical analysis indicates that this limitation may be attributed to their tendency to converge to sharp minima, which are known to correlate with weaker generalization performance. To this end, we propose a new method called Sassha that stably minimizes sharpness within the framework of second-order optimization. Sassha converges to flat solutions and achieves state-of-the-art performance within this class. Sassha also performs competitively to widely-used first-order, adaptive, and sharpness-aware methods. Sassha achieves this efficiently through lazy Hessian updates, to which it is robust, and does so without requiring extra hyperparameter tuning. Moreover, Sassha exhibits strong resilience to label noise. All of these are rigorously assessed with extensive experiments.
Nonetheless, there are still many limitations to be addressed to further improve this work. Some examples may include, but are not limited to, extending experiments to various models and different data of extreme scales, as well as developing theoretical properties such as convergence rate, generalization bound, and implicit bias, all to more rigorously confirm the value of Sassha. Seeing it as an exciting opportunity, we plan to investigate further in future work.
Acknowledgement
This work was partly supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (IITP-2019-0-01906, Artificial Intelligence Graduate School Program (POSTECH), RS-2024-00338140, Development of learning and utilization technology to reflect sustainability of generative language models and up-to-dateness over time), and the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) ( NRF-2022R1F1A1064569, RS-2023-00210466, RS-2023-00265444).
Impact statement
This paper contributes to advancing second-order optimization, with potential implications for both theoretical insights and practical applications. While our work does not present immediate concerns warranting specific emphasis, we recognize that progress in this field may have broader societal impacts. We remain committed to engaging in discussions on the broader implications of our research should the need arise in the future.
@@ Appendix A Sharpness Measurements for Other Settings
Table 7: Sharpness measurements of the solutions found by seven different optimizers and their generalization on CIFAR-10/100 and Wikitext-2. Approximate second-order methods tend to yield highly sharp solutions and poor generalization compared to SGD; Sassha and M-Sassha effectively recover this. Here, we measure sharpness in terms of maximum Hessian eigenvalue $\lambda_{max}(H)$, trace of Hessian $\operatorname{tr}(H)$, gradient-direction sharpness $\delta L_{\text{grad}}$, and average sharpness $\delta L_{\text{avg}}$, along with generalization using validation loss $L_{val}$ and accuracy $\text{Acc}_{val}$.
@@
$\lambda_{max}(H)$ $\operatorname{tr}(H)_{\times 10^{3}}$ $\delta L_{\text{grad}}$ $\delta L_{\text{avg}\times 10^{-3}}$ $L_{\text{val}}$ $\text{Acc}_{\text{val}}$
@@
ResNet20 SGD 107±4.370 1.380±0.010 0.840±0.304 0.690±0.390 0.295±0.008 92.03±0.32
SAM 58±2.980 0.730±0.040 0.171±0.038 0.461±0.240 0.119±0.002 92.85±0.07
AdaHessian 23048±29932 189.5±240.6 4.538±1.634 198.7±266.0 0.260±0.006 92.00±0.17
Sophia-H 3606±303.0 31.24±2.628 6.120±1.634 18.11±1.000 0.316±0.002 91.81±0.27
Shampoo 647066±419964 3900±1825 166.3±48.00 2177189±1628993 0.381±0.028 88.55±0.83
M-Sassha 129±17.00 1.580±0.080 1.551±0.684 1.025±0.360 0.234±0.003 92.36±0.23
Sassha 78±5.090 0.860±0.030 0.184±0.053 0.388±0.704 0.209±0.001 92.98±0.05
ResNet32 SGD 56±5.100 0.800±0.040 0.560±0.219 0.196±0.146 0.309±0.002 92.69±0.06
SAM 45±2.670 0.580±0.020 0.107±0.005 0.753±0.351 0.128±0.001 93.89±0.13
AdaHessian 1746±1018 17.06±10.24 4.599±1.710 5.518±3.623 0.278±0.006 92.48±0.15
Sophia-H 7167±2755 18.82±5.500 9.399±2.283 7.915±3.397 0.394±0.010 91.99±0.08
Shampoo 717553±93129 4523±629.7 162.1±123.2 105322±82246 0.348±0.008 90.23±0.24
M-Sassha 283±10.00 3.960±0.100 2.986±1.133 1.300±0.969 0.211±0.010 93.18±0.30
Sassha 47±1.880 0.590±0.020 0.136±0.019 0.714±0.090 0.177±0.002 94.09±0.24